OpenStudy (anonymous):

a cuboid with a square base is to be made with 200 cm^2 of material. Show that the height, h=(50/x)-(x/2)

OpenStudy (anonymous):

@hartnn please give me some hints like last time ! :)

OpenStudy (anonymous):

let the base side be x then S.A=2(lb+bh+lh) 200=2(x^2+hx+hx) result follows

OpenStudy (anonymous):

@ksw19372

OpenStudy (anonymous):

still confused..

hartnn (hartnn):

S.A=2(lb+bh+lh) got this ?

hartnn (hartnn):

or which part of @Galaxyy 's explanation did u not get ?

OpenStudy (anonymous):

i know its really basic but i dont get what i and b i know h is for height..

OpenStudy (anonymous):

l-length b-bredth h-height

OpenStudy (anonymous):

now its easier :) thank you to both of you !

OpenStudy (anonymous):

ur welcome